Salaries for senior Civil engineers per year have known mean of 71 K dollars and standard deviation 3.9 K dollars. Provided that the salaries per year for senior Civil engineers have a nearly symmetric b ell-curve distribution, what percen % of Civil engineers have salary between 671 K and 71 K dollars? Enter a mathematical Points possible: 5 Unlimited attempts. Score on last attempt: 0. Score in gradebook: 0 Message instructor about this question more Licenseplease help me do this
Answer #1

Solution:

Civil engineers have salary between 67.1 and 71 k dollars are 34.13 %

Step 1: Sketch the curve The probability that 67.1<X<71is equal to the blue area under the curve 0.12 0.10 0.08 0.06 0.04 0.02 0.00 67.1 0.02 58 62 70 72 78 80 82 Step 2 Since μ-71 and σ 3.9 we have 67.1-71 X-μ 71 P(67.1 < X < 71 ) = P (67.1-71 < X-μ<71-71 ) = P Since Z-Z-μ 3.01 and -71 67 d71 1 0 we have 3.9 P(07.1 < X < 71 ) = P (-1 < Z < 0 ) Step 3: Use the standard normal table to conclude that P(-1< Z <0) 0.3413